### v6

# Tensor polarization and the dissipative damping of vector meson in QCD Medium

Abstract
Unexpectedly large and puzzling spin alignment, and thus tensor polarization, of vector mesons has been observed in heavy-ion collisions. Given that tensor polarization represents a fluctuation of spin, we derive, for the first time, a fluctuation-dissipation relation for tensor polarization, where the polarization is proportional to first-order hydrodynamic gradients (e.g., the shear-stress tensor), with dissipative coefficients depending on the particle’s damping properties, as characterized by its spectral function. Employing relativistic hydrodynamics at finite density, we find that dissipative contributions can generate substantial spin alignment. We provide illustrative examples (by tuning one coefficients $\alpha_{{\rm sh}}$) that generate a beam energy, $p_{T}$, and centrality dependence of spin alignment resembling those observed in experiments, offering insights into these puzzling phenomena and demonstrating its potential as a “spectrometer” for in-medium vector mesons.
Introduction.— Spin has played a central role in advancing our understanding of the microscopic quantum world since its discovery. Due to its quantum nature, the measurement of spin is inherently non-deterministic, and we only obtain statistical distributions. Therefore, statistical quantities such as the mean value of the spin vector $\langle S_{i}\rangle$(with $i=x,y,z$) and spin fluctuations $\langle S_{i}S_{j}\rangle$ are essential for characterizing a particle’s spin properties. For spin-1/2 particles, the mean value of the spin vector $\langle S_{i}\rangle$ is sufficient to determine the complete spin information, i.e., the spin density matrix, of the particle. However, for spin-1 particles, in addition to the vector spin, the symmetric and traceless part of the fluctuation tensor $\langle S_{i}S_{j}\rangle$, specifically $\langle S_{(i}S_{j)}-\bm{S}^{2}/3\rangle$, carries an additional 5 degrees of freedom to characterize the spin density matrix. This quantity, with proper normalization, is referred to as the tensor polarization of spin-1 particles, and has been investigated in many fields in physics Bourrely et al. (1980); Haftel et al. (1980); Holt et al. (1981); Schulze et al. (1984); Abbott et al. (2000); Chen et al. (2016, 2020); Gregory et al. (2017).
In heavy-ion collisions, besides the widely studied vector polarization in experiments Adamczyk et al. (2017a); Adam et al. (2018); Niida (2019); Adam et al. (2019); Acharya et al. (2022) and theory Liang and Wang (2005a); Becattini et al. (2013); Karpenko and Becattini (2017); Li et al. (2017); Sun and Ko (2017); Pang et al. (2016); Liu et al. (2020); Liu and Yin (2021); Fu et al. (2021); Becattini et al. (2021a, b); Fu et al. (2022); Wu et al. (2022a), the tensor polarization, related to spin fluctuations, of vector mesons has also been proposed to be an observable Liang and Wang (2005b). More concretely, the proposed observable is called spin alignment, which refers to the deviation of the “00” component of the spin density matrix from its equilibrium limit: $\delta\rho_{00}=\rho_{00}-1/3$. This quantity is directly proportional to the $zz$-component of the tensor polarization: $\delta\rho_{00}\propto\langle 2S_{z}^{2}-S_{x}^{2}-S_{y}^{2}\rangle/3=T_{zz}$(quantize along $\hat{z}$), characterizing how fluctuations in the system are unevenly distributed between the transverse xyx-y plane and the longitudinal $z$-direction. However, recent measurements in experiments  Abdallah et al. (2023); Abelev et al. (2008); Acharya et al. (2020); Singha (2021); ALI (2022) have discovered that the value of the spin alignment is orders of magnitude larger than typical theoretical expectations Becattini and Piccinini (2008); Becattini et al. (2008); Xia et al. (2021); Becattini (2022), accompanied by puzzling behaviors.
Efforts have also been made from various perspectives Sheng et al. (2020); Müller and Yang (2021); Weickgenannt et al. (2022); Sheng et al. (2022, 2023) to understand this puzzle. However, the origin of this unexpected spin alignment remains an open question Becattini (2022). These studies have primarily established non-dissipative relations between spin alignment and other quantities, such as vorticity Becattini and Piccinini (2008); Becattini et al. (2008); Xia et al. (2021); Becattini (2022) or strong electromagnetic fields Sheng et al. (2020, 2022, 2023), with the latter being proposed as an explanation in experimental studies Abdallah et al. (2023). However, following Einstein’s discovery of the physics of Brownian motion, it is understood that the fluctuation of quantities can be naturally linked to dissipative processes. Given that spin alignment represents an anisotropic fluctuation of spin, it is natural to ask whether a fluctuation-dissipation relation can be established in this context—a question that has yet to be explored in the study of spin alignment. Therefore, in this work, we present two key findings from our investigation:
(1) Theoretically, we establish, for the first time, a fluctuation-dissipation relation for tensor polarization, where novel dissipative contributions are discovered.
(2) Phenomenologically, the new dissipative effects exhibit promising features that shed light on this puzzling spin alignment observed in experiments.
Fluctuation-dissipation relation .— To further illustrate the fluctuation nature of the tensor polarization and spin alignment, we begin with a simplified example of the spin density matrix $\rho_{ss^{\prime}}$ (or $\rho$) in the particle’s rest frame, where the tensor polarization density $T_{ij}$ can be expressed in terms of the standard $3\times 3$ spin operators $S^{i}$ (standard representation Li and Liu (2025)) as $\mathcal{T}_{ij}\propto-\text{Tr}\{\rho(S_{(i}S_{j)}-\bm{S}^{2}/{3})\}3n$. Correspondingly, the spin alignment can be expressed as $\delta\rho_{00}\propto T^{zz}\propto\text{Tr}\{\rho(S_{x}^{2}+S_{y}^{2}-2S_{z}^{2})/3\}$. As shown, both $T_{\mu\nu}$ and $\delta\rho_{00}$ can be interpreted as the fluctuation of spin.
To obtain useful formulations for these quantities, the theory is developed in the following context: vector mesons with momentum $\tilde{p}^{\mu}=(\varepsilon_{\bm{p}},\bm{p})$ at density $n(\varepsilon_{u})$ propagate in a locally equilibrated hydrodynamic medium, where their tensor polarization ${\cal T}^{\mu\nu}$ can be characterized by hydrodynamic variables such as inverse temperature $\beta=1/T$, the flow velocity $u^{\mu}(x)$, symmetric flow gradients $\xi_{\lambda\gamma}\equiv\beta^{-1}\partial_{(\lambda}(\beta u)_{\gamma)}$, etc. To study general polarization phenomena in this context, we have developed a comprehensive theoretical framework “Zubarev response approach”, which is detailed in Supplementary Materials (Supp) and in our long theory paper Li and Liu (2025). Especially, in the long paper, we illustrate how our “Zubarev response approach”, rooted rigorously in thermal field theory, can systematically calculate various polarization phenomena, including the dissipative and non-dissipative contribution on equal footing, and lay out the path to incorporating non-perturbative methods, such as lattice methods and the functional renormalization group.
There are many new dissipative and non-dissipative effects predicted by the “Zubarev response approach” to the spin alignment problem, where the fluctuation and dissipation relation are naturally encoded. In this letter, we will focus on discussing one of the main discoveries–shear-induced tensor polarization effects (SITP), using it as an example to highlight how dissipative contributions, overlooked by the community for decades, can impact the phenomenology of spin alignment physics. Using SITP effect as an example, the uneven spin fluctuation characterized by tensor polarization ${\cal T}^{\mu\nu}$ can be expressed as
${\cal T}^{\mu\nu}=\tilde{\Delta}^{\langle\mu}_{\lambda}\tilde{\Delta}^{\nu\rangle}_{\gamma}[\beta n(\varepsilon_{u})\alpha_{{\rm sh}}\xi^{\gamma\lambda}]$ (1)
which naturally relates to the dissipative processes in-medium through a T-odd dissipative coefficient as
$\alpha_{\rm sh}$ $=\frac{4\varepsilon_{\bm{p}}\pi}{\beta n(\varepsilon_{\bm{p}})}\int^{\infty}_{0}d\omega\frac{\partial n(\omega)}{\partial\omega}(\omega^{2}-\varepsilon_{\bm{p}}^{2})A^{2}(\omega,\bm{p})$
$\approx-\frac{2\Delta\varepsilon_{\bm{p}}}{\Gamma_{\bm{p}}}+2\frac{\Delta\varepsilon_{\bm{p}}}{\Gamma_{\bm{p}}}\frac{\Delta\varepsilon_{\bm{p}}}{T}+\frac{\Gamma_{\bm{p}}}{2T}\,$, (2)
where the spectral function $A(\omega,\bm{p})$ controls the dissipative damping of a particle’s propagation as $D(t,x)\propto\int d\omega\,e^{-i\omega t+\bm{p}\cdot x}A(\omega,\bm{p})$. The projectors in Eq. (1) are defined as $\tilde{\Delta}^{\mu\nu}=-\eta^{\mu\nu}+\tilde{p}^{\mu}\tilde{p}^{\nu}/m^{2}\;$ and $\tilde{\Delta}^{\langle\mu}_{\;\lambda}\tilde{\Delta}^{\nu\rangle}_{\;\gamma}=\tilde{\Delta}^{(\mu}_{\;\lambda}\tilde{\Delta}^{\nu)}_{\;\gamma}-\tilde{\Delta}^{\mu\nu}\tilde{\Delta}_{\lambda\gamma}/3$. The second line of Eq.(Tensor polarization and the dissipative damping of vector meson in QCD Medium) employs a quasi-particle approximation, where the $\Delta\varepsilon_{\bm{p}}$ is the in-medium energy shift and is T-even. The $\Gamma_{\bm{p}}$ is the width of the particle, i.e. twice damping rate, of the particles in-medium, which is the T-odd and is the source of dissipation in this calculation. As shown, the $\alpha_{\rm sh}$ is an odd power of $\Gamma_{\bm{p}}$ , so it is T-odd and dissipative.
It is worth noting that terms like $\Gamma_{\bm{p}}/(2T)$ do not have mass suppression, allowing them to survive in non-relativistic limits. Consequently, this tensor polarization, especially the SITP effect, has a potential to be observed in low-energy physics experiments, such as plasma or cold atom systems. Meanwhile, directly extracting the in-medium mass shift and width from the shape of the invariant mass spectra of kaon pairs Pal et al. (2002) or dileptons Li and Ko (1995) is highly challenging. The observed spectra in experiments Abdallah et al. (2023) are dominated by the spin-aligned $\phi$ mesons that survive to kinetic freeze-out with a vacuum spectral function. As shown in Eq. (1) and Eq. (Tensor polarization and the dissipative damping of vector meson in QCD Medium), these in-medium spectral properties are naturally related to the transport coefficients generating the spin alignment, allowing us to use the spin alignment in reverse as a probe to the spectral properties of the QCD matter. In the following, we will use simple setups for the width and mass shift to study how the SITP effect manifests in phenomenology, which reversely illustrates this point.
Phenomenological setup.— We adopt the commonly used freezeout assumption for spin physics Becattini et al. (2019); Fu et al. (2021); Becattini et al. (2021b). Specifically, we consider a physical scenario in which the $\phi$ meson, carrying an equilibrated spin alignment, forms as early as the late stages of the QGP Shuryak and Zahed (2004); Liu and Rapp (2018). Depending on conditions such as centrality, beam energy, etc., the fireball evolution varies, leading the spin alignment to effectively freezeout at the QGP phase, mixed phase, or hadronic phase. This contributes to the observed centrality and beam energy dependence (see latter). After this “spin freeze-out,” the spin alignment stays approximately unchanged. The $\phi$ mesons surviving through the kinetic freeze-out will recover the vacuum spectral function but retain a “memory” of spin alignment from the spin freeze-out. Their decay products no longer undergo re-scattering and will carry all this information to the final observables.
Figure 1: Compare to experimental data Abdallah et al. (2023) by tuning one parameter $\alpha_{\rm sh}(p,T)$($\alpha_{\rm sh}(\Delta m,\Gamma)$), using the same experimental cuts: (1) $\alpha_{\rm sh}$ for different cases and their $\rho_{00}$; (2) Beam energy dependence; (3) $p_{T}$ dependence; (4) Centrality dependence.
We employ the standard Cooper-Fry-like formula to characterize the freezeout of spin alignment, where the $\delta\rho_{00}$ can be expressed with tensor polarization ${\cal T}^{\mu\nu}$ as
$\delta\rho_{00}(\hat{n}_{\text{pr}},\bm{p})=\frac{\int d\Sigma^{\lambda}p_{\lambda}\,{\cal T}^{\mu\nu}(x,\bm{p})\hat{n}_{\mu}(\bm{p})\hat{n}_{\nu}(\bm{p})}{d\Sigma^{\lambda}p_{\lambda}3n(\varepsilon_{u})}\,.$ (3)
The $\Sigma^{\lambda}$ is the hyper-surface satisfying the freeze-out condition, and $\hat{n}^{\mu}$ is a shorthand for $\epsilon^{\mu}_{s=0}$, which is related to a three-dimensional unit vector $\hat{n}_{\text{pr}}$ by a Lorentz boost, i.e., $\hat{n}^{\mu}=[\Lambda(\mathbf{p}).\hat{n}_{\text{pr}}]^{\mu}$, with $\hat{n}_{\text{pr}}$ being the polarization vector $\epsilon^{\mu}_{s=0,\text{pr}}$ in the particle rest frame. The $\hat{n}_{\text{pr}}$ is taken out/in-plane ($\hat{y}/\hat{x}$) direction in this work. We utilize the CLVisc Pang et al. (2016) at finite density Wu et al. (2022b) for the hydrodynamic simulation at various beam energies and centralities. Unless otherwise stated, the settings are directly inherited from the work Wu et al. (2022b); Chen et al. (2024) without re-tuning in our event-average qualitative study. In the following, we will evaluate the spin alignment with $\alpha_{\rm sh}$ from two setups, “physically motivated coefficients” or microscopic “quark-meson (QM) model coefficients”, aiming to extract model-independent features of SITP effects.
Comparison with data and “spectrometer” .— In this section, we tune only one parameter, $\alpha_{\text{sh}}(p,T)$, for the SITP effect and compare the resulting spin alignment with experimental data. As previously explained, $\alpha_{\text{sh}}(p,T)$ is a function of $\Delta m$ and $\Gamma$. Therefore, we will examine how sensitively the data can, in turn, serve as a “spectrometer” for vector meson properties.
We first construct a physically motivated $\alpha_{\rm sh}$ with the help of Eq. (Tensor polarization and the dissipative damping of vector meson in QCD Medium). Motivated by Ref Rapp (2001); van Hees and Rapp (2008); Vujanovic et al. (2009); Liu and Rapp (2020, 2018) and the QM model discussed later, we introduce some typical momentum dependence using a Lorentz contraction $\gamma_{\bm{p}}^{-1}=m/\varepsilon_{\bm{p}}\equiv\sqrt{1-v_{\bm{p}}^{2}}$ factor for the width and its corresponding in-medium mass shifts as $\Gamma_{\bm{p}}=\Gamma_{0}/\gamma_{\bm{p}}$ and $\Delta\varepsilon_{\bm{p}}=\Delta m/\gamma_{\bm{p}}$, The $\Gamma_{0}$ and $\Delta m$ are chosen to be $\Gamma_{0}=0.1$ GeV and $\Delta m=\{-0.03,0.01,0.05\}$ GeV, where $\alpha_{\rm sh}$’s for these three cases are shown in upper panel of the $1^{\text{st}}$ figure in Fig. 1. The $\Gamma_{0}=0.1$ GeV is not unreasonable large for the particles close to the phase boundary in heavy-ion collisions since even for heavy mesons like $J/\psi$ can have a width around 0.1 GeV in this region He et al. (2022); Liu and Rapp (2018). For lighter vector mesons, such as $\phi$, it is reasonable to expect a width of this order. As shown, depending on the sign and magnitude of a small mass shift (ranging from 30-30 to $50$ MeV), $\alpha_{\rm sh}$ can reach values as large as “1” and as small as “0.8-0.8”. The mass shifts around the phase boundary can be huge (few hundreds MeV) Liu and Rapp (2018, 2020). The sign-flipping behavior is mainly due to the competition between the $2\Delta\varepsilon_{\bm{p}}/\Gamma_{\bm{p}}$ and $\Gamma_{\bm{p}}/2T$ in Eq. (Tensor polarization and the dissipative damping of vector meson in QCD Medium).
With these coefficients, the spin alignments $\rho_{00}$ generated within the framework are shown in the lower panel of the first figure of Fig. 1. The magnitude of spin alignment roughly ranges from -2% to the 4%, which is orders of magnitude larger than previous predictions Xia et al. (2021) and close to what is observed in experiments Abdallah et al. (2023). Meanwhile, for all three $\Delta m$ cases, the $p_{T}$ dependence scales with the $p$ dependence of the coefficients. In particular, for the $\Delta m=0.01$ GeV case, the $p_{T}$-dependent sign-flipping behaviors observed in experiments are reproduced.
As a cross check of hydrodynamic calculation, we also estimate the $\rho_{00}$ with $\bm{p}=0$ in the medium rest frame using some simple math following Eq. (1):
$\delta\rho_{00}\propto\frac{1}{9}\alpha_{\text{sh}}\beta(\partial_{z}u^{z}+\partial_{x}u^{x}-2\partial_{y}u^{y})$
$\delta\rho_{00}=\frac{1}{9}\alpha_{\text{sh}}\beta\partial_{z}u^{z}=\frac{1}{9}\alpha_{\text{sh}}\frac{\beta}{\tau},\text{Ideal Bjorken}\,.$ (4)
With a positive $\alpha_{\text{sh}}$ and the fact that usually $\partial_{z}u^{z}>\partial_{x}u^{x}>\partial_{y}u^{y}$, we will get positive spin alignment. Using an ideal Bjorken hydrodynamic solution with $\alpha_{\text{sh}}\sim 1$ and $\tau\sim 6$ fm, we will have $\rho_{00}\sim 6\%$, which is not very far from the 4% obtained from a realistic hydrodynamics simulation shown in Fig. 1.
By correlating beam energies and centralities with spectral properties, such as mass shift or width, we can generate spin alignments with rich behavior. In Fig. 1, the $2^{\text{nd}}$ figure demonstrates that a change of less than 40 MeV (ansatzes in figures) in the in-medium $\phi$ mass with beam energy mimics the observed beam energy dependence of $\rho_{00}$ for $\phi$ and $K^{*}$. The out-plane $\rho_{00}$ is automatically separated from the in-plane one. Although the splitting size depends on the details of the hydrodynamic simulation, the general hierarchy is robust and resembles that observed in experiments. We employ different illustrative ansatzes for the spectral properties of $K^{*}$ and $\phi$, reflecting their different in-medium interactions and freezeout stages: $K^{*}$ freezes out in the late hadronic phase (positive mass-shifts Ilner et al. (2017)), while $\phi$ is assumed to freeze out earlier (QGP or early hadronic phase).The corresponding $p_{T}$ dependence of $\rho_{00}$ is shown in the $3^{\text{rd}}$ figure, which can accommodate a $p_{T}$-dependent sign-flipping behavior. Correlating spectral properties with $\sqrt[3]{dN/d\eta}$, fitted linearly with centrality Adare et al. (2016) ($4^{\text{th}}$ figure, more in Fig. 6 Supp), can generate a sign-flipping $\rho_{00}$ dependence, mimicking those observed in experiments.
Microscopically, the spectral function’s beam energy dependence may relate to the chemical potential, which increases at lower energies. The physics correlating centrality with spectral properties is more subtle. In the next section, we use a microscopic quark-meson model to illustrate this physics further. It is worth noting that there are still many other potential effects, such as the strong field effect Sheng et al. (2023), initial/local fluctuations  Kumar et al. (2023); Fang and Pu (2025), etc. Kumar et al. (2024); Zhang et al. (2025); Li (2024), that might be entangled. Nevertheless, our calculations support the importance of the thermal-fluctuation-dissipation contribution and demonstrate that spin alignment is a candidate to serve as a sensitive “spectrometer”, showing that a few-tens-MeV change in mass-shift and width can generate distinguishable behaviors in experimental observables.
Quark-meson model coefficients.— To further justify physically motivated coefficients and obtain the model-independent insights, we have done a microscopic calculation on the coefficients and used it to calculate the spin alignment in this section. We first evaluate the coefficients $\alpha_{\text{sh}}$ numerically using Eq. (Tensor polarization and the dissipative damping of vector meson in QCD Medium) for $\phi$ meson with in-medium spectral functions obtained via a 1-loop quark meson model as discussed in Supp (similar ideas in Ref. Dong et al. (2023)). With coupling $g_{\phi}=2$, these coefficients are plotted in Fig. 2 for the cases with the effective strange quark masses $m_{s}=\{0.3,0.42,0.5\}$ GeV, which are relevant mass scales of strangeness at the QGP phase, mix phase, and hadronic phase. In the hadronic phase, this mass represents a effective in-medium kaon mass (see discussion later).
Figure 2: (1) The $\alpha_{\rm sh}$ coefficients for different cases; (2) $p_{T}$-dependent $\rho_{00}$ for ’green’ coefficients in (1); (3) $p_{T}$-integrated $\rho_{00}$ as a function of centrality; (4) illustration of $m_{s}$ change with centrality and freezeout $T$.
As shown in Fig. 2, microscopically calculated $\alpha_{\text{sh}}$ generates a $p$ dependence and sign-flipping behaviors similar to those plotted in Fig. 1, and they share similar physics but there are some extra physics of the blue curves. For $m_{s}=0.5$ GeV, which is an effective kaon mass, the $\phi$ meson mass is so close to the $2m_{s}$ threshold that $\Gamma_{\bm{p}}$ becomes much smaller, making $\Gamma_{\bm{p}}/(2T)<2\Delta\varepsilon_{\bm{p}}/\Gamma_{\bm{p}}$ and $\alpha_{\text{sh}}$ negative for all momenta. This threshold physics will be similar if we switch to a hadronic Lagrangian, as illustrated in our following work Sun et al. (2025), where $\alpha_{\text{sh}}$ can be $(-1.5\sim-0.5)$ using a chiral perturbative hadronic Lagrangian or a face value of $\alpha_{\text{sh}}\sim-2$ from Ref Haglin and Gale (1994). Meanwhile, we note that the quark meson model used here does not generate large positive $\alpha_{\rm sh}$. This is because the current medium effects are dominated by radiative effects, and the effective mass shift is positive, making it difficult to generate large positive values of $\alpha_{\rm sh}$. However, in other situations where collision broadening effects dominate, large widths and negative mass shifts might occur simultaneously Liu and Rapp (2018), leading to larger positive $\alpha_{\rm sh}$. Therefore, in a more realistic situation, we can expect a large positive coefficient similar to those shown in Fig. 1 in certain scenarios.
The $2^{\rm nd}$ and $3^{\rm rd}$ panels in Fig. 2 illustrate the $p_{T}$ dependence and the centrality dependence of the spin alignment from the QM model. For these figures, we actually include all new effects (splitting, zeroth order) discussed in Supp and we confirm that the SITP effect is the dominant one in this setup. As shown, microscopic and more complete calculation can recover the sign-flipping $p_{T}$ and centrality dependence illustrated Fig. 1. While the $p_{T}$ dependence has already been discussed, the“centrality-freezeout correlation” is illustrated in the $4^{\rm th}$ panel that we utilize hydrodynamic simulations with a different (spin) freezeout temperature $T_{\rm fz}$ correlated with centrality, which is then corrected with $m_{s}(T_{\rm fz})$ and those $\alpha_{\rm sh}$ coefficients.
The plotted “centrality-freezeout correlation”, where $T_{fz}$ increases with centrality (more peripheral), resembles the trend in kinetic freeze-out Adamczyk et al. (2017b). The hadronic phase lifetime is largely linear with $\sqrt[3]{dN/d\eta}$ Knospe et al. (2016, 2021); Adam et al. (2017); Acharya et al. (2021). Since $\sqrt[3]{dN/d\eta}$ also decreases linearly with centrality (see Fig.  6), the lifetime drops from 7 fm (0%) to 2 fm (60%) at top RHIC. At large centrality, the shorter hadronic phase, which inhibits re-equilibration in a later, lower-temperature stage, effectively raises the “freezeout temperature”. Meanwhile, the mass-dropping scheme employed for the effective s-quark (or other constituents, such as kaons) is mostly motivated by chiral symmetry restoration physics Buballa (2005); Li (2016) and emerging low-lying QGP collective modes Liu and Rapp (2018) that might be hundreds of MeV below the condensate mass. On the other hand, for those who have concerns about the idea of the strange-masses-melting scenario, we can go back to the Fig. 1, where the mass-shifts and width changes required is less than 100 MeV. This is not a strong requirement or unreasonable assumption. Collisional interactions in the QGP phase can generate negative mass shifts as large as a few hundred MeV Liu and Rapp (2018), while radiative ($\phi\rightarrow\bar{K}K\,\text{or}\,\bar{s}s$ type) interactions generate positive mass-shifts as shown in our QM model or Haglin and Gale (1994). Therefore, if the interaction type transitions from collisional to radiative type as it crosses the phase boundary, we can expect a sign-flipping in mass shift, leading to sign-flipping in spin alignment.
Besides published data, preliminary experimental data Wilks (2025) suggest a strong rapidity dependence at lower energies. While the SITP effect is largely boost-invariant at mid-rapidity, this dependence can be accounted for by including the zeroth-order term with coefficient $\alpha_{0}$ (see Eq. (I.1) Supp). Tuning both $\alpha_{0}$ and $\alpha_{\rm sh}$ can accommodate this new feature without qualitatively altering the features shown in Fig. 1. A large rapidity dependence would imply a $\phi$ meson with a longitudinal/transverse mass splitting at the order of 100 MeV (see large splitting in other particles Yeo et al. (2024)), a possible case for a strongly-coupled medium (see Fig. 5 Supp).
Summary.— We have developed a Zubarev response approach based on thermal field theory and utilized it to study the tensor polarization and spin alignment of vector mesons, leading to two key theoretical findings:
(1) We derive for the first time a fluctuation-dissipation relation for tensor polarization and spin alignment, revealing spin alignment is naturally related to damping properties and, consequently, to the spectral functions.
(2) We discover a novel shear-induced tensor polarization (SITP) effect at first order of hydrodynamic gradients, which can generate significant spin alignment.
Phenomenologically, by tuning the parameter $\alpha_{\rm sh}$, the SITP effect can generate spin alignment that depends on beam energy, centrality, $p_{T}$, quantization axes, and particle type ($\phi$, $K^{*}$), resembling those observed in experiments, including various sign-flipping behaviors. Further including zeroth-order effect with an extra parameter, $\alpha_{0}$, can produce a large rapidity dependence.
In the context of the symmetry-based gradient expansion (see Eq. (I.1)), this study demonstrates that terms proportional to $\partial^{(\mu}\xi^{\nu)}$ and $u^{\mu}u^{\nu}$ already provide strong explanatory power for current experimental observations. This offers a promising path toward understanding spin alignment physics within the general framework of the relativistic hydrodynamic model, often referred to as the “standard model” in heavy-ion physics.
Meanwhile, we also demonstrate that a few-tens-MeV change in mass-shifts and width can significantly change the transport coefficients associated with the above gradients, generating distinguishable behaviors in experimental observables and making spin alignment serve as a promising “spectrometer” candidate for vector mesons.
To fully realize this “spectrometer” blueprint, we need to investigate additional contributions beyond the leading orders to refine the relation between the spectral function and spin alignment. This effort also requires more sophisticated studies of in-medium spectral properties, alongside further improved phenomenology and simulations. Investigating other non-equilibrium effects and novel contributions proposed in Refs. Sheng et al. (2023); Müller and Yang (2021) is also necessary to understand their quantitative impact. Finally, as a separate point, the newly discovered SITP effect is robust against variations in interaction details and is universal in both relativistic and non-relativistic cases, making it interesting to investigate experimentally in low-energy physics, such as plasma or cold atom physics Gregory et al. (2017).
Acknowledgements.
Acknowledgments We gratefully acknowledge the valuable contribution from Yi Yin, who definitely deserves a key position in the author list, but waive the authorship generously. We extend special thanks to Xiangyu Wu for sharing the most recent CLVisc code and providing helpful guidance for our simulations at various beam energies in later versions. We are also especially grateful to Longgang Pang for providing hydrodynamic profiles for our early numerical studies. We also thank Youyu Li, Xinxiang Liu, and Tan Luo for engaging in valuable discussions, digitizing experimental plots, and providing guidance on simulations. We thank Hengtong Ding, Xiaojian Du, Aiqiang Guo, Koichi Hattori, Min He, Che-Ming Ko, Yutie Liang, Shu Lin, Rapp Ralf, Kaijia Sun, Subhash Singha, Shuzhe Shi, Yifeng Sun, Aihong Tang, Biaogang Wu, Dilun Yang, Wenbin Zhao for valuable discussions. SL is supported by NSFC No. 12205090. Fundamental Research Funds for the central Universities. FL is supported by NSFC No. 12105129.